### v2

# The Superstatistics of confined ants: From serenity to panic

Abstract
Nonequilibrium systems with strong parameter fluctuations are difficult to describe with ordinary Statistical Mechanics. Superstatistics –that can be seen as statistics of an statistics– has emerged as a tool able to describe these complex systems. It has been successfully applied, for example, to hydrodynamic turbulence, internal convection, and even DNA architecture, but not to macroscopic biological systems. Here we present fully controlled laboratory experiments in confined ants both in normal and in “panic” conditions. Using a data-driven superstatistical model, we reproduce key signatures of their motion: from non-Gaussian velocity distributions driven by a Log-normal diffusion coefficient, to clustering away from the source of danger information.
Superstatistics; Langevin; ants; panic; collective motion
I Significance Statement
We introduce Superstatistics as a study tool in the animal world, and report its Log-normal variant in a living system. Our study suggests that “panicked” ants, when exposed to a threat, can transmit information to their nestmates through a barrier they cannot physically cross. We explain both normal and panic dynamics, identifying non-Gaussian velocity densities, time scale separation, and the statistics of the diffusion coefficient. We also reveal distinct behavioral regimes in the ants’ collective panic response and how it relates with the individual ant memory and cluster formation. Furthermore, our findings indicate that optical signals or simple antennation are not significant mechanisms for panic transmission. We believe that these discoveries provide a foundation to uncover the biological origin of Log-normal type diffusion in confined environments.
II Introduction
Nature displays a diversity of collective biological behaviors ranging from the microscopic to the macroscopic scales, and from high confined environments to the open space: the movement of cells during wound heal Vedel et al. (2013); Angelini et al. (2011); Zhang et al. (2010) and the flocking of birds Cavagna et al. (2010); Ballerini et al. (2008); Pearce et al. (2014) illustrate quite well those scenarios.
Here, we focus on the collective motion of confined macroscopic individuals, which is a common scenario in eusocial organisms, such as ants, who often live in high density societies. Leaf-cutter ants like Atta Insularis build complex nests made of networks of tunnels and chambers within which they live, move, take care of eggs, and even harvest fungus. While there is an extensive body of work that qualitatively describes the activity of ants in both open and confined spaces, its detailed quantitative understanding is still challenging Hölldobler and Wilson (1990); Gordon (2010); Sumpter (2006).
In terms of quantitative modelling, one may start invoking the Langevin equation, which captures the finding that animals tend to maintain their direction of motion during some characteristic (persistence) time Méndez et al. (2016); Viswanathan et al. (2011); Gurarie et al. (2017); Santana-Filho et al. (2020). Namely,
$\frac{d}{dt}v_{i}(t)=-\gamma v_{i}(t)+\sigma\xi_{i}(t)$ (1)
where $v_{i}$ denotes the $i$-th component of velocity; $\gamma$, a positive parameter that encloses directional persistence; and $\sigma$, the intensity of a Gaussian white noise, $\xi_{i}$. A more complete overview of the statistical properties of Equation (1) was made in 1931 and it is known as Ornstein-Uhlenbeck process Uhlenbeck and Ornstein (1930), nowadays a reference in ecological models for persistent movement Méndez et al. (2016).
In addition to this, a superposition of different dynamics on different time scales often describes properties of driven nonequilibrium systems Beck (2001), e.g., the fast jiggly dynamics of a Brownian particle subject to slow environmental fluctuations of a given diffusivity Wang et al. (2009); Chechkin et al. (2017). This is known as Superstatistics (which stands for superposition of statistics), and one of its key merits is the ability to explain the emergence of non-Gaussian densities Beck and Cohen (2003); Metzler (2020).
The theory is connected to Langevin systems and, within the framework of Statistical Mechanics, can be interpreted by considering a system that is composed of small regions or cells that are temporarily in local equilibrium, showing spatio-temporal fluctuations of some intensive quantity, such as the inverse temperature $\beta=2\gamma/\sigma^{2}$ in Equation (1). It is assumed that the system reaches local equilibrium on a timescale $\tau_{1}$ much smaller than the time $\tau_{2}$ needed for $\beta$ to vary among cells. Consequently, within each region, $\beta$ could be considered approximately constant. Then, during time intervals of the order of $\tau_{2}$ and inside each cell of a system following the dynamics of Equation (1), local Gaussian velocity stationary distributions of the form:
$p(v_{i}|\beta)=\left(\frac{\beta}{2\pi}\right)^{1/2}e^{\displaystyle{-\beta v_{i}^{2}/2}}$ (2)
arise. This means that to describe the long-time behavior, the marginal distribution for the $i$-th component of velocity $v_{i}$,
[…]
should be computed. Here, $f(\beta)$ is the probability density function of $\beta$, and a relation such as Equation (3) explains the emergence of many non-Gaussian behaviors reported in the literature Queirós (2008); Rizzo et al. (2004); Porporato et al. (2006); Briggs and Beck (2007); Chen and Beck (2008).
By performing careful experiments, here we use the behavior of confined ants in normal and panic conditions as a case study for superstatistical analysis. We quantify the motion of ants that are confined into a squared, quasi-2D cell. A number of interesting features arise: from non-Gaussian ant velocity distribution heavily influenced by speed-dependent noise intensity, to the existence of a fluctuating diffusivity, characteristic of non-Gaussian diffusion emergence in single-particle tracking experiments performed in soft and biological matter Wang et al. (2012); Chen and Beck (2008). Then, we expose the confined ants to the effect of danger information: ants in a neighbouring cell that are chemically excited transmit “panic information” to the individuals under study through a specially-designed barrier that cannot be physically crossed. A collective output is built within minutes: in spite of their short individual memory Gordon (2010), the ants form a cluster whose center shifts away from the barrier through which the danger information has emerged. In other words, the individual short-term memory has become a long-term collective memory when danger information is turned on. We propose a superstatistical model able to quantitatively reproduce all the experimentally observed features, both before and after panic. To the best of our knowledge, this is the first documented occurrence of Superstatistics in the animal world and Log-normal Superstatistics in any living system.
III Experimental details
A system consisting in two 3D-printed cells, each of $6\times 6\times 1\,\mathrm{cm^{3}}$ and separated by a barrier, was designed to investigate the possible transmission of danger information by ants from one cell to another (see Figure 1). Ants are placed in both cells and filmed during $5$ minutes. Then, an insect repellent is added at the center of the left cell and another $10$ minutes of film is taken. The left and right cells will be referred to as DANGER and SAFE, respectively. In a CONTROL experiment the same procedure is followed, except no ants are introduced into the DANGER arena. We hypothesize that, if ants in the SAFE arena move away from the barrier after the addition of the repellent a significantly larger distance than they do during a CONTROL experiment, danger information has been transmitted by the “panicked” ants in the DANGER cell to the ants in the SAFE cell beyond any “passive” diffusion of the repellent through the barrier.
Figure 1: Experimental setup. (A) Upper panel: General view of the device. Lower panel: sketch of the three types of barriers used for the protocols GLASS, THICK and THIN. The latter is also used for the CONTROL protocol (see text). (B) Top view of ants near a thick and thin barriers. Short distance contacts are not allowed by the thick barrier. (C) Lateral view of an ant near a thin barrier. As a reference, the length of the ants bodies is around $1\,\mathrm{cm}$.
In an attempt to isolate the possible mechanisms of danger information transmission, various types of barriers were employed (lower panel of Figure 1A). The glass barrier consists of a transparent glass wall aimed at filtering optical signals as the primary mechanism for the transfer of panic information. The thick barrier is made in such a way that ants on both sides can make contact with their extremities (e.g., their respective antennas) but cannot antennate the bodies of individuals in the opposite arena. In contrast, the thin barrier allows such contacts (see Figure 1C). We term our protocols GLASS, THICK and THIN, corresponding to each type of barrier.
A total of $10$ repetitions for each protocol was made and ants were used only once in order to avoid any effects associated to previous individual experience Grüter and Czaczkes (2019). The experiments were performed in workers of the ant species Atta insularis, called bibijagua in Cuba, that have been extensively studied both in natural and laboratory conditions Altshuler et al. (2005); Nicolis et al. (2013); Noda et al. (2006); Tejera et al. (2016); Serrano-Muñoz et al. (2019); Reyes et al. (2016) (see Figure S1 in Supplementary Information for a high resolution picture of a typical individual).
IV Collective motion: Tracking the Center of Mass
Using image processing techniques, we tracked the center of mass (CM) in both cells for each protocol. Data was processed using yupi, a Python library to handle (tracking, transform, generate, store, analyze, visualize) trajectory data Reyes et al. (2023). As illustrated in Figure 2A-C for typical experiments, the motion of the CM at the DANGER cell is always near its geometrical center, independently from the protocol used. A similar situation holds for the SAFE cell, with the conspicuous exception of THIN, where the CM tends to move away from the barrier after the repellent is added into the DANGER arena. Figure 2E-H shows the time evolution of the $x$-coordinate of the CM at the SAFE cell with respect to its starting position, corresponding to an average between the repetitions of each type of experiment.
Figure 2: Center of mass. Snapshots of experiments (A) GLASS, (B) THICK, (C) THIN and (D) CONTROL (see text for details). Green, blue and red traces correspond to the motion of the center of mass for consecutive $5$-minutes time windows from the start of an experiment (the snapshots were taken at the end of each video). (E-H) Time evolution of the center of mass measured from its starting position along the horizontal direction (i.e., $x-x_{0}$) inside the SAFE (right) cell for each protocol ($x=0$ is taken at the center of the cell). Light-colored shadows indicate standard deviation. Notice that a repellent is added in the DANGER (left) cell at $t=5$ minutes into the repellent reservoir seen as a white spot in (D).
The net displacement of the CM, defined as the difference between its final and initial $x$-coordinate averaged over the $15$th and $5$th minutes, was computed (see Figure S2 on Supplementary Information). A non-parametric Mann-Whitney test Mann and Whitney (1947) was used to test the hypothesis that the pair of samples THIN-GLASS, THIN-THICK and THIN-CONTROL come from populations with identical means. We find that the three p-values are $0.0017$, $0.0004$ and $0.0373$, respectively. Hence, we argue that only in the THIN protocol the addition of repellent induces an “escape response” in the ants trapped in the SAFE cell.
With this we are also saying that the CONTROL experiment shows that the mere diffusion of the repellent from DANGER to SAFE does not produce a significant shift of the CM in the SAFE cell. So, the escape response is mediated by some kind of information transmitted by panicked ants from DANGER to SAFE. (A video containing an example of each kind of experiment, including the evolution of the CM, can be found in Movie S1 as Supplementary Information).
Figure 2C,G also reveals an interesting behavior of the ants in the SAFE arena as they escape from the barrier. Immediately following the repellent injection, the velocity of the CM rapidly spikes, indicating a fast dynamic response that lasts for about one minute. This is followed by a slower process in which the ants form a stable cluster at a certain distance from the barrier, a retreat that is quite distinct from the immediate reaction seen when ants are directly exposed to insect repellent Altshuler et al. (2005). Further details of this behavior will be discussed later.
V Motion analysis before panic
Figure 3: Suppressing boundary effects. (A) Acceleration as a function of position. The inset plot shows acceleration absolute value as a function of the distance from the walls. A characteristic length $\ell_{w}=0.6\,\mathrm{cm}$ is defined as the effective wall thickness. (B) The grey sections have been removed from a sample trajectory, as they are within a distance $\ell_{w}$ from the walls: in order to ensure that statistics is not contaminated by boundary effects, we will only use for future analysis the non-gray sections.
Having established that the THIN protocol promotes the propagation of danger information through the barrier, we studied individual trajectories of ants confined in the SAFE cell only using thin barriers. We used the CONTROL experiments to establish a “baseline” model of how ants behave when confined in the SAFE cell, considering only their interactions with each other and with the cell walls. We chose CONTROL protocol because, in this setup, ant interactions with the left wall (i.e., the barrier) are influenced solely by the barrier itself, without interference with the ants in the DANGER arena. Then, in the next section, we will build upon this model and extend it to the case where ants also interact with those in the DANGER cell and get danger information.
Data consists in approximately $20$ randomly selected ants that were tracked manually during different repetitions of the CONTROL protocol. Trajectories were taken at intervals of $1/3$ of a second, so position samples are spaced at intervals of $\Delta t\approx 0.3\,\mathrm{s}$.
Figure 4: Experimental results before panic. (A) Mean (blue and orange) and fluctuation intensity (dark and light green) of parallel and perpendicular acceleration components for a given speed value. The mean parallel acceleration scatters linearly while perpendicular components remains at zero. Both components of standard deviation are equal and speed-dependent. The fitting line is an exponential of the form of Equation (4b). Although they are plotted on the same axes, note that $\mu(v)$ and $\sigma(v)$ have different units (i.e., $\mathrm{cm\,s}^{-2}$ and $\mathrm{cm\,s}^{-3/2}$, respectively). (B) Velocity autocorrelation function (VACF) up to a lag time of $2\,\mathrm{s}$. The exponential decay is governed by a persistence time $\gamma^{-1}=0.53\,\mathrm{s}$. (C) Average of the local flatness during different time intervals (see Equation (5)). The dashed line indicates the flatness of the Gaussian. In the inset plot, a typical realization of the diffusion coefficient, $D(t)$ (dark red), shows how it changes slowly compared to $v_{x}(t)$ (dark blue). Both quantities are plotted as a function of time over approximately 50 seconds, though neither time nor value ticks are displayed for clarity. (D) Probability density of the diffusion coefficient $D$ (in units of its standard deviation) extracted from velocity time series as indicated in Equation (6). The data was found to fit a Log-normal distribution. For a better visualization of the tail fit, a logarithmic plot is shown as an inset up to $15$ standard deviations.
To eliminate boundary effects, we measured first the extent of wall “forces”. Hence, experimental acceleration components (i.e., $a_{x}$ and $a_{y}$) before repellent injection, were arranged based on increasing values of $x$- and $y$- position components, respectively. Both acceleration time series were binned and a running average was taken (see Figure 3A). The ants slow down as they approach the walls, revealing the influence of the boundaries as if they were confined by a soft, square-box-like potential. Consequently, a plot of acceleration absolute value against the distance from the wall was made, and a characteristic length $\ell_{w}=0.6\,\mathrm{cm}$ resulted from fitting to an exponential function (inset of Figure 3A). Henceforth, the analysis in the rest of this section is conducted with the sections of all trajectories that were consistently at least $\ell_{w}$ units away from every boundary, to discard boundary effects.
In order to look for the first insights of a theory that explains ant behavior before panic was induced, we computed the expected value of acceleration and its fluctuations against instantaneous speed by defining parallel and perpendicular components to the moving direction. That is, defining the unit vectors, mutually orthogonal, $\mathbf{e}^{(\|)}(t)$ and $\mathbf{e}^{(\bot)}(t)$, parallel and perpendicular to the direction of velocity $\mathbf{v}(t)$, respectively as
$\mathbf{e}^{(\|)}=\mathbf{v}/v\quad\text{and}\quad\mathbf{e}^{(\bot)}=(-e^{(\|)}_{y},e^{(\|)}_{x})^{\top}$
we obtain a new coordinate system that changes direction with $\mathbf{v}(t)$, where we projected the instantaneous acceleration. Note that $v\equiv|\mathbf{v}|$.
Figure 4A shows that the acceleration parallel to the velocity at any time $t$ is a random vector whose average is proportional to $-v$, while orthogonal accelerations locally average to zero for all speed values. The standard deviation of both components, however, are not independent of velocity, as demonstrated in the upper part of the same figure, and within their error bars, they exhibit the same non-linear dependence on $v$. Therefore, for the mean of the parallel component, $\langle\mathbf{a}^{(\|)}\rangle$, and the intensity of both the parallel and perpendicular fluctuations, $\langle(\mathbf{a}^{(\|,\bot)}-\langle\mathbf{a}^{(\|,\bot)}\rangle)^{2}\rangle^{1/2}\Delta t^{1/2}$, which we denote as $\mu(v)$ and $\sigma(v)$, respectively, we propose the following relations:
$\mu(v)$ $=-\gamma v$ (4a)
$\sigma(v)$ $=\Sigma+\Delta\Sigma\,(1-e^{\displaystyle{-v/v_{\sigma}}})$ (4b)
where $\gamma^{-1}$ is a persistence time. The first term, $\Sigma$, accounts for the intensity of acceleration fluctuations at rest, while the second term describes its increase with the velocity, mediated by a characteristic speed $v_{\sigma}$ and a proportionality factor, $\Delta\Sigma$. Note that, in order to obtain $\sigma(v)$, the standard deviation of acceleration time series was scaled by $\Delta t^{1/2}$ to ensure dimensional consistency with a continuous stochastic process, where the white noise has units of $\mathrm{s}^{-1/2}$. We also tested other models for the function $\sigma(v)$, such as general power laws of the form $a+b\,v^{c}$, but none performed better than the exponential model in Equation (4b), as they could, for instance, produce unphysical values for $\sigma(v=0)$. Equation (4b) was fitted and the numerical values of the parameters are reported in Table 1. These results strongly suggest to begin with a Langevin equation such as Equation (1) as a model for ant movement, with a slight modification that allows for $\sigma$ to be expressed as $\sigma(v)$, a function of speed.
Table 1: Values for model parameters fitted to data.
$\ell_{w}$ $a_{0}$ $\gamma^{-1}$ $\Sigma$ $\Delta\Sigma$ $v_{\sigma}$
$\mathrm{(cm)}$ $\mathrm{(cm/s^{2})}$ $\mathrm{(s)}$ $\mathrm{(cm\,s^{-3/2})}$ $\mathrm{(cm\,s^{-3/2})}$ $\mathrm{(cm/s)}$
0.6 1.0 0.53 0.34 1.99 1.99
$s$ $\tau_{f}$ $\tau_{s}$ $L$ $t_{p}$ $\tilde{x}$
(adim) $\mathrm{(s)}$ $\mathrm{(min)}$ $\mathrm{(cm)}$ $\mathrm{(min)}$ $\mathrm{(cm)}$
1.27 20 3.2 6.0 5.5 3.8
Alongside, the temporal correlation of ant velocities decay exponentially, as shown in a plot of the velocity autocorrelation function (VACF) depicted in Figure 4B. With this, we found a persistence time of $\gamma^{-1}=0.53\,\mathrm{s}$ (also reported in Table 1), which is consistent with the inverse of the slope of parallel acceleration plot shown in Figure 4A. Let’s now look in more detail into the stochastic part of accelerations.
Figure 5A shows the probability density function (PDF) for the zero-mean parallel and perpendicular acceleration components, i.e., $\mathbf{a}^{(\|,\bot)}-\langle\mathbf{a}^{(\|,\bot)}\rangle$, in units of its standard deviation. The non-Gaussian shape highlights the need for a model with non-Gaussian noise. A straightforward assumption is to allow the noise intensity to be stochastic. This leads us to the domain of Superstatistics, as the more basic consideration following a constant noise intensity is that it should be a slowly varying process. Note that in Figure S3, the noise autocorrelation function (NACF) appears as a flat function of time close to zero, so acceleration fluctuations are not correlated at the time scales under study.
A dynamical reason for the non-Gaussian behavior seen in Figure 5A has been proven to be due to Superstatistics whenever a superposition of local Gaussians (2) whose variance fluctuates on a large spatio-temporal scale arise. Therefore, it is needed to check if time series does contain separate time scales. For that, as suggested by Beck et al. (2005), a time average of the local flatness, $\kappa_{i}$, of velocity components, $v_{i}$, was computed for increasing time intervals, $\tau$:
[…]
Here, $T$ is the overall measurement time and $\langle v_{i}\rangle$ is the average of the $i$-th velocity component on a time window, $\tau$. The result is shown in Figure 4C, where a dashed line was drawn to indicate the flatness of the Gaussian. The time $\tau_{2}$ at which this line intercepts $\kappa_{i}(\tau)$ corresponds to a time scale $\tau_{2}=3.3\,\mathrm{s}$. Having pointed out that the time scale at which velocity fluctuates is $\tau_{1}\equiv\gamma^{-1}=0.53\,\mathrm{s}$, the ratio $\tau_{2}/\tau_{1}\approx 6$ implies that there exist some intensive parameter, $\beta(t)$, that fluctuates on a temporal scale 6 times larger than the one of the velocity, $\mathbf{v}(t)$.
The slowly varying stochastic process $\beta(t)$ can then be determined by noticing that $\beta^{-1}$ is the variance of the local Gaussians (2). Namely,
$\beta(t)=\frac{1}{\langle v_{i}^{2}(t)\rangle_{\tau_{2}}-\langle v_{i}(t)\rangle^{2}_{\tau_{2}}}$ (6)
where $\langle\cdot\rangle_{\tau_{2}}$ means that the average is taken over a time window of the order of $\tau_{2}$. So, we looked for a suitable power $p$ such that $\beta^{p}$ could make the dynamics fall into one of the universality superstatistical classes, e.g., $\chi^{2}$ or Log-normal Beck (2006). We found that, indeed, when $p=-1$, it is the diffusion coefficient, $D\propto\beta^{-1}$, the slowly varying intensive parameter. Furthermore, Figure 4D shows that it follows a Log-normal density that, in its standardized form, looks like:
$f(D)=\frac{1}{D\sqrt{2\pi s^{2}}}\exp{\left(-\frac{\log^{2}D}{2s^{2}}\right)}$ (7)
and whose shape parameter, $s$, is given in Table 1. Note that the fit was performed within a range of up to $15$ standard deviations (inset of Figure 4D). Caution must be exercised when fitting experimental data as Log-normal Superstatistics and stretched exponentials since they can produce similar fits and typical experimental conditions often do not reach the asymptotic tails Beck (2006). In the inset plot of Figure 4C typical realizations of $v_{x}(t)$ and $D(t)$ are depicted, and the time scale separation can be qualitatively verified by observing how one process changes faster than the other.
Figure 5: Statistical distributions before panic: experiment and theory. Dots stand for experimental data, where light/dark blue and orange represent parallel-perpendicular/$x$-$y$ components. Dark green lines are predictions of the superstatistical model (8) (in short, SSModel). (A) Probability density function (PDF) of parallel and perpendicular components of acceleration fluctuation about its mean value. The non-Gaussian density is explained by the model as the product of a Log-normal with a Gaussian. (B) Probability density function of $x$ and $y$ velocity components. Three classical densities, i.e., Gaussian, Laplace and Power law (Lévy stable), are fitted and shown as dashed lines. A high peak at low velocities is predicted by our model as well as tails heavier than a Gaussian.
Based on these results, and without introducing free parameters, we used the persistence time $\gamma^{-1}$, the speed-dependent noise intensity $\sigma(v)$, and the slowly varying diffusion coefficient $D(t)$ –which follows a Log-normal distribution with $s$ as the shape parameter– to develop a superstatistical Langevin equation to model the velocity component $v_{i}(t)$ of ants in the SAFE cell prior to the induction of danger, taking the form:
$\frac{d}{dt}v_{i}(t)=-\gamma v_{i}(t)+\sigma(v)\sqrt{2\gamma^{2}D(t)}\,\xi_{i}(t)$ (8)
Here, at any time instant $t$, the diffusion coefficient $D(t)$ and the Gaussian $\xi_{i}(t)$ are used to generate the stochastic variable $\eta(t)=\sqrt{2\gamma^{2}D(t)}\,\xi_{i}(t)$. Then $\eta(t)$ is standardized by dividing by its standard deviation to be finally scaled by $\sigma(v)$. A Runge-Kutta method of strong order $1.5$ within a Milstein scheme to compute Stratonovich-type integrals was used for numerical integration Burrage and Burrage (1996).
The PDF of parallel and perpendicular acceleration fluctuations can be successfully explained by Equation (8), as shown in Figure 5A. The model captures the non-Gaussian essence of the stochastic part that characterizes changes in velocity. Furthermore, Figure 5B displays experimental probability densities for the $x$ and $y$ velocity components. Apart from fluctuations due to the finite nature of experimental data, no discrepancies are observed between the shapes of the distributions, which is consistent with the spatial homogeneity and isotropy of the surface of the arena. The model also predicts the observed peak at low velocities and tails that extend far from a Gaussian yet bellow a power law, in line with the fact that accelerations spread around their mean value with standard deviations depending on $v$ along with the existence of the superstatistical Log-normal diffusion coefficient.
VI Panic response
Figure 6: Experimental results and model prediction after panic. (A) Local flatness as a function of lag time for different time span until the end of the experiments. The plots gather sections of the trajectories in the time intervals $[0,5)\,\text{min},[5,6)\,\text{min},\dots,[14,15)\,\text{min}$. In the right corner, the final time of each interval is displayed. Time scale separation is preserved under panic. (B) Probability density function (PDF) of centered parallel and perpendicular components of acceleration fluctuation. The diffusion coefficient retains its Log-normal statistics so the shapes of the acceleration distributions are the same. (C) Density of cartesian velocity components. The histograms become short-tailed functions after panic induction. (D) Temporal and speed dependence of the intensity of acceleration fluctuations for the parallel (blue) and perpendicular (orange) components after panic was induced. Continuous lines show fits to functions of the form (4b). (E) Temporal dependence of the parameters of the Equation (4b) that fit the functions in (D). Two of them exhibit a sudden change at minute 9.2, and only the noise intensity at rest, $\Sigma(t)$, seems to remain unaltered. (F) Net displacement of the center of mass for the THIN protocol (same curve of Figure 2G) and prediction of the model (9).
We now discuss the next 10 minutes after the repellent was injected into the DANGER cell, using data from approximately 50 trajectories of ants in the SAFE arena that were randomly chosen and tracked across all repetitions of the THIN protocol.
Firstly, we found that the mean acceleration maintains the same linear relationship with speed values (Figure S4(left)). A clear time-scale separation is still present, with the slow time scale, $\tau_{2}$, remaining approximately constant over time (Figure 6A). Note that just after a duration of $\tau_{2}$ units, the local kurtosis spreads out, allowing us to assume local equilibrium within the panic regime whenever $t<\tau_{2}$. No significant changes were observed in the persistence time $\gamma^{-1}$ (Figure S4(right)), or in the statistics of the diffusion coefficient, $D$, as indicated by the shape invariance of centered acceleration densities in Figure 6B. The tails of the velocity distributions become stretched, as shown in Figure 6C. This observation aligns with the gradual clustering dynamics under panic conditions, as depicted in the SAFE cell of Figure 2C. Non-gaussianity again arises as a direct combination of a $v$-dependent noise intensity and Superstatistics.
Moreover, the intensity of the parallel and perpendicular fluctuations of acceleration components, $\sigma(v)$, was the only observable that appeared to be significantly affected following the onset of panic. Figure 6D shows the time evolution of $\sigma(v)$ along with the fitting according to Equation (4b). Both components still remain indistinguishable over time. To further analyze their time-dependent behavior, we inspect the evolution of the parameters in Equation (4b). Figure 6E shows that the three parameters may be assumed nearly constant until around the 9th minute of the experiment, when two of them, $\Delta\Sigma(t)$ and $v_{\sigma}(t)$, increase suddenly and quickly relaxes back to its initial value, while $\Sigma(t)$ stays invariant. For later simulations, $\Delta\Sigma(t)$ and $v_{\sigma}(t)$ were fit to contiguous back-to-back exponential functions with matching growth and relaxation rates, selected for their simplicity in characterizing rapid changes. The timing at which this peak occurs will subsequently be used as a key measure of one of the most significant timescales in the ants’ mechanism in response to panic.
As mentioned previously, the collective behavior of ants in the SAFE cell is mediated by a fast dynamics that span approximately one minute as soon as the repellent is injected. Before that, ants were dispersed throughout the arena. Therefore, when panic is induced in the 5th minute of the experiment, it can be assumed that an ant at the barrier in SAFE is immediately informed and becomes “activated”. After a certain characteristic time, $\tau_{f}$, this ant communicates the danger information to another nestmate, resulting in two activated ants. This cascade continues, and it ends after $T=1\,\mathrm{min}$ when the total number of individuals, $N_{T}=20$, in the arena are informed. Consequently, the number of activated ants at time $t$, $N(t)$, obeys the differential equation $\dot{N}=\lambda_{f}N$ with the initial condition $N(0)=1$, which implies that $\tau_{f}=T/\ln N_{T}=20\,\mathrm{s}$.
While it is true that between minutes 5 and 6 more than one ant reaches the barrier, the process would then be mediated by multiple ants seeking information, retaining it, and triggering a cascade. This, along with the finite nature of information retention, leads us to conclude that a more simplified mechanism explaining the shift in the CM can be understood as a single “effective ant” representing several informed ants. From the moment this effective ant reaches the barrier, it remains activated and triggers a panic avalanche among its nestmates.
When all ants are informed in the 6th minute of the experiment, a slower dynamics arises. It can be characterized by the time between the moment when all members were activated and the peak observed in Figure 6E at minute $9.2$, i.e., by a characteristic time $\tau_{s}=3.2\,\mathrm{min}$. Hence, an extended model based on (8) that incorporates boundary effects and panic dynamics mediated by the times $\tau_{f}$ and $\tau_{s}$, was evaluated as follows:
$\begin{split}\frac{d}{dt}v_{i}(t)&=-\gamma(v_{i}(t)-\tilde{v}_{i}(t))\\ &\quad\,+\sigma(v,t)\sqrt{2\gamma^{2}D(t)}\,\xi_{i}(t)+F_{w}(r_{i})\end{split}$ (9a)
$\sigma(v,t)$ $=\Sigma(t)-\Delta\Sigma(t)\,e^{\displaystyle{-v/v_{\sigma}(t)}}$ (9b)
$F_{w}(r_{i})$ $=a_{0}(e^{\displaystyle{-r_{i}/\ell_{w}}}-e^{\displaystyle{-(r_{i}-L)/\ell_{w}}})$ (9c)
[…]
Here, $r_{i}$ stands for the $x$-$y$ components of the ants’ position. $F_{w}(r_{i})$ in (9c), with units of acceleration, mimics the wall forces as shown in Figure 3A (main plot) with intensity $a_{0}=1.0\,\mathrm{cm/s^{2}}$ resulted from the fitted line shown in the inset plot, and $L=6.0\,\mathrm{cm}$ being the length of the arena.
On the other hand, the drift velocity $\tilde{\mathbf{v}}(t)$ in (9d) is a vector with a null $y$ component since the repellent effect causes asymmetries only in the direction perpendicular to the barrier. In contrast, the $x$ component accounts for some drift dominated by the fast and slow time scales, $\tau_{f}$ and $\tau_{s}$, respectively. The function is an asymmetric sigmoid centered at the midpoint of the fast dynamics (i.e., at $t_{p}=5.5\,\mathrm{min}$) and saturates at a certain velocity $F_{w}(\tilde{x})/\gamma$. This implies that there is a balance of forces for large $t$: the force to the left from the right wall and the “panic force” that continues pushing to the right. This is why the drift velocity has been expressed in terms of the position $\tilde{x}$, which is the only free parameter of the model and was fixed at the distance where the ant cluster remains at rest at the end of the experiment.
With this, we have quantitatively reproduced the experimental motion of the CM using the superstatistical model with parameters directly measured from the experiments. Figure 6F presents the $x$-coordinate of the CM from its initial position (as shown earlier in Figure 2G) alongside the prediction of the extended model (9), plotted as a dark green solid line. The numerical values for all model parameters are provided in Table 1.
In addition, it is worth noting that not only does the diffusion coefficient fluctuates on the same time scale before and after panic is induced, but its Log-normal statistics also remains invariant. This leads to the idea that the emergence of a Log-normal type diffusion might be a fundamental feature in the interaction of ants under confinement, and thus it “survives” even in states of danger, which is revealed in the robustness of the PDFs of the acceleration over the whole length of the experiments, shown in Figure 6B.
Similarly, the fact that $\Sigma(t)$ remains nearly constant after the repellent is injected indicates that panic does not significantly alter the noise intensity at rest. Furthermore, the total noise intensity, $\Sigma+\Delta\Sigma$, appears to be “remembered”, as $\Delta\Sigma(t)$ returns to its pre-peak value after its sharp rise, maintaining the overall noise level.
It is instructive to condense the superstatistical treatment into a narrative of the ants response to panic in the SAFE cell, as time goes by. Immediately after the injection of the repellent in the DANGER cell, there is a period from minute 5 to minute 6 that can be identified in Figure 6F as the onset rising of the blue curve from zero. It corresponds to a “cascade dynamics” where individual ants get danger information from nestmates in the DANGER cell through the barrier, and pass it to other ants in the SAFE cell within a characteristic time $\tau_{f}=20\,\mathrm{s}$. Interestingly, this number is of the order of the “ant memory” time reported in the literature Gordon (2010). A second stage occurs from minute 6 to minute $9.2$ approximately, where the blue curve in Figure 6F increases steeply, with a characteristic time $\tau_{s}=3.2\,\mathrm{min}$. Within that stage, informed ants “rectify” their fluctuating motion in the direction away to the barrier, thus provoking a rightward shift of the CM. A final stage takes place from minute 9.2 to minute 15, where the ants cluster stabilizes, and the CM reaches its final position due to the saturation of the $\tilde{\mathbf{v}}(t)$ “force”. This results from the balance between the rightward “panic repulsion” from the barrier due to the input of danger information, and the leftward “mechanical repulsion” due to the presence of the right wall of the right cell.
Finally, there is a remaining puzzle that we have not discussed yet: the time stability of the cluster, in spite of the short “individual memory” of each ant. A possible explanation is suggested by the fact that our clusters are not absolutely static: individual ants eventually separate from the group, explore, and return. We hypothesize that the information feedback provided by those ants from the boundaries of the right cell is rapidly transmitted to the rest of the densely-packed ants, so the cluster-formation behavior is constantly reinforced. This can be better understood by inspecting Figure S5, which shows the decrease in the probability of excursions to the barrier. Given that the left edge of the cluster is qualitatively observed at a distance no less than approximately $2\,\mathrm{cm}$ from the barrier, Figure S5 depicts the probability of finding an ant at a distance less than or equal to $2\,\mathrm{cm}$ from the barrier, as a function of time. After the repellent is introduced, barrier exploration decreases, though it never completely reaches zero. So, the establishment of the cluster within the last 5 minutes of the experiments illustrated in Figure 6F, is associated to the analogous portion of Figure S5, where a small (but steady) ant traffic between the cluster and the barrier is minimally “refreshing” the panic information.
VII Conclusions
We have presented controlled laboratory experiments with confined leaf-cutter ants of the species Atta insularis, where ants in a confined space were studied in “normal” conditions for 5 minutes, and in the presence of panic information emerging from the left side of the cell, for another 10 minutes. The experiments show a shift of the center of mass of the ants towards the right as panic is switched on, and the formation of a stable cluster in the last minutes of the experiment.
We observed non-Gaussian probability density functions for both the ants’ acceleration fluctuations and velocities, a slowly varying diffusion coefficient and distinct behavioral regimes in the ants’ collective response to panic. Furthermore, our findings indicate that optical signals or simple antennation do not play a significant role in this panic transmission, raising the possibility of more complex communication mechanisms at work.
We were able to quantitatively reproduce all the above observations using a superstatistical, data-driven model involving a robust Log-normal type diffusion, with all relevant parameters extracted from experimental data. The model suggests that, when panic is triggered, individual ants gather information at the source of danger and transmit it to others within a time scale expected for the memory of an individual. This leads to a cascade of events, which turn on a collective motion away from the source of panic and the emergence of a stable cluster over a larger characteristic time. Future research may focus on identifying which biological hidden “degrees of freedom” –perhaps corresponding to the precise ant-ant communication mechanism– potentially acts in a cascade-like fashion to give rise to the Log-normal diffusion.
VIII Acknowledgements
A. Reyes gratefully acknowledges the Abdus Salam ICTP STEP program for partial financial support of this research. Antonio Celani, Deborah Gordon, Roberto Mulet and Tarek Tohme contributed with enlightening discussions and comments. We thank Antonio Serrano for helping in the design of Figure 1A. E. Altshuler gratefully acknowledges an Invited Professorship supported by the CNRS at the University of Lyon.